OpenStudy (anonymous):

Find the value of A(10). round your answer to the nearest hundredth

OpenStudy (anonymous):

what is A(10)

OpenStudy (anonymous):

algebra 1 common core\[A10\]

OpenStudy (anonymous):

value?

OpenStudy (anonymous):

not getting

OpenStudy (anonymous):

\[1/(s ^{2 + \sqrt{2}} s + 1)\]

OpenStudy (anonymous):

i have to show my work

OpenStudy (anonymous):

can you show me work i have radiation in a hour and have litttle time thanks

OpenStudy (anonymous):

i hav ans but i want explanation how they reached from this step to that step

OpenStudy (anonymous):

\[\sqrt{2} e^{-t/\sqrt{2}} \sin (t/\sqrt{2})\] this is ans but i want explanation of middle steps shortly

OpenStudy (anonymous):

it just says find value of A(10) round to nearest hundreths